Karen is 15 years older than Kyla. 13 years from now the sum of their ages will be 109. Give the current ages of Karen and Kyla.
Answer by checkley71(8403)

kyle's age=x
karen's age=x+15
in 13 years they will be
kyle=(x+13)
karen=(x+15+13)
so adding their ages 13 years from now they will=109
(x+13)+(x+15+13)=109
2x+41=109
2x=109-41
2x=68
x=68/2
x=34 years old is kyle's age now.
proof
(34+13)+(34+15+13)=109
47+62=109
109=109